OpenStudy (anonymous):

(f 0 g) (10) f(x)= 2x+3 g(x)= x-1/3

OpenStudy (anonymous):

f(g(10)) let 'g' become your new 'x', we'll call this 's' so now s=x-1/3 f(g(x)) = 2s+3 ( because s is your new 'x') now f(g(x))=(fog(x)) = 2(x-1/3))+3 expand and then sub x=10 into it

OpenStudy (anonymous):

Do you want the answer or entire solution?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

Not to be a party pooper but the website isn't for giving solutions, it;s to aid your understanding. Helping you understand the method and get to the answer yourself is fine, but giving you the answer isn't really what the site is for, it won't help you get it.

OpenStudy (anonymous):

I got the answer